A parallelogram has sides of length 1.3 feet and 3.2 feet. What is the perimeter?
1
Expert's answer
2019-09-25T09:48:07-0400

The perimeter of parallelogram = 2(side A + side B )


= 2 (1.3 + 3.2) = 2(4.5) =9 feet